Find the equation of the line passing through points (2, 5) and (-4, -7).


1
Expert's answer
2021-09-06T19:16:20-0400

Slope of the line: "m=\frac{5+7}{2+4}=2."

Equation of the line: "y-5=2(x-2)" or "y=2x+1."